# Mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium: An effective-dipole analysis

### v1

@@ Abstract
The mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium have been investigated, in terms of hypothetical effective dipoles that depict how the positive and negative charge in each particle are separated. Our findings, which revealed that it is possible for polarization-induced opposite-charge repulsion to occur at short interparticle separations if the dielectric constant of the medium is greater than the dielectric constants of both spheres, provide insights into the physics of charge separation behind such counterintuitive behaviour.
@@ I Introduction
While the classic Coulomb’s law Jackson (2021) describes the electrostatic force between two point charges in vacuum, it does not necessarily describe the electrostatic force between two charged dielectric particles. In some cases, there could be a significant deviation from the Coulomb force at small separations, resulting in the occurrence of polarization-driven like-charge attraction Feng (2000); Bichoutskaia et al. (2010); Xu (2013); Freed (2014); Qin et al. (2016, 2016); Lindgren et al. (2018); Qin (2019); Chan (2020); Li et al. (2024); Gorman et al. (2024). On the other hand, charged particles immersed in an electrolyte could also exhibit anti-Coulomb behaviour, including cases of like-charge attraction  Allahyarov et al. (1998); Levin et al. (1999); Linse and Lobaskin (1999); Moreira and Netz (2001); Nagornyak et al. (2009); Zhao et al. (2016); Buyukdagli (2017); Dos Santos and Levin (2019); Wills et al. (2024) and also cases of opposite-charge repulsion Aranda-Espinoza et al. (1999); Zhang and Shklovskii (2004); Leunissen et al. (2005); Wills et al. (2024), but such behaviour is not purely driven by the polarizability of the particles nor that of the surrounding medium. While it is physically impossible for two dielectric particles in vacuum to exhibit opposite-charge repulsion, by taking into account the polarizability of the medium it would be worthwhile to investigate whether opposite-charge repulsion could also occur as a pure consequence of polarization effects, as in the case of like-charge attraction between two dielectric particles in vacuum.
@@
To understand the physical mechanisms behind the occurrence of polarization-driven like-charge attraction between two polarizable particles, much research has been conducted by means of 1) accurate evaluations of the interparticle force or energy Bichoutskaia et al. (2010); Khachatourian et al. (2014); Lindgren et al. (2016, 2018); Lian and Qin (2018); Li et al. (2024) that take into account all possible higher-order interactions, or 2) effective-dipole-based modelsChan (2020) or computationsLindgren et al. (2018); Li et al. (2024) that provide a physical picture of charge separation in each particle. For the former, a multiple scattering formalism Freed (2014); Qin et al. (2016, 2016); Qin (2019) has been developed, where analytic solutions for generally many-body systems could be obtained through an application of the perturbative many-body expansion to the interaction energy. On the other hand, there have been various attempts to develop efficient numerical methods for accurate computations of the force or energy: conventional approaches, such as the finite element method Feng (2000) and the boundary element method Gorman et al. (2024); Ruan et al. (2022); Hassan and Stamm (2021); Bramas et al. (2021), typically involve the discretization of a volume or surface. At small interparticle separations, such approaches become computationally expensive, because a high resolution of the discretization is required. For this reason, some more efficient approaches have been developed. These include 1) the image charge method  Wang and Wang (2013); Xu (2013) where, for each particle, the polarization field is approximated as a field generated by the image charges in the other particle and by the multiply-reflected image charges due to the presence of two interfaces, 2) the method of moments  Bichoutskaia et al. (2010); Derbenev et al. (2016); Lindgren et al. (2016); Siryk et al. (2021) where the polarization-related potential is described in terms of a multipole expansion, the coefficients of which are determined by substituting the expansion into the Poisson equation and the corresponding boundary conditions, and 3) a hybrid computational method that employs the image charge method to compute the fast varying near-field interactions and the method of moments to solve the remaining smooth far-field interactions.
@@
In this research, we used a combination of the hybrid computational method with the effective-dipole approach to investigate the possibility of polarization-induced opposite-charge repulsion between two dielectric spheres in close proximity inside a polarizable medium. Our findings revealed that such a possibility exists if the dielectric constant of the medium is greater than the dielectric constants of both spheres, in which case the repulsive interaction corresponds to an anti-parallel alignment of the particles’ effective dipoles Lindgren et al. (2018) and hence a dominant like-charge repulsion between the polarization charges of the particles.
@@
This paper is organized as follows. In Sec. II, we present a theoretical model of two charged dielectric spheres, on which our research is based, along with a brief review of the hybrid computational method Gan et al. (2019) that we employed to compute the interparticle force and the effective-dipole moment of each particle  Lindgren et al. (2018). In Sec. III, we present our findings on the nature of interparticle interactions and on the corresponding orientations and relative positions of effective dipoles for four different scenarios: (i) Like-charge repulsion or attraction, where the dielectric constant of the medium is smaller than the dielectric constants of both particles; (ii) Opposite-charge attraction (no repulsion), where the dielectric constant of the medium is smaller than the dielectric constants of both particles; (iii) Like-charge repulsion (no attraction), where the dielectric constant of the medium is larger than the dielectric constants of both particles; (iv) Opposite-charge attraction or repulsion, where the dielectric constant of the medium is larger than the dielectric constants of both particles. In Sec. IV, we summarize our findings and discuss their implications.
@@ II.1 Model of dielectric spheres in a polarizable medium
Figure 1: Schematic illustration of two equal-sized dielectric spheres, immersed in a medium of dielectric constant $k_{0}$ and separated by a centre-to-centre separation $h$. For sphere $i\in\left\{1,2\right\}$ with a radius $a$ and a dielectric constant $k_{i}$, its free charge $q_{i}$ is concentrated at the particle’s centre $\mathbf{z}_{i}$, and its bound charge spans the particle’s surface.
Consider two equal-sized dielectric spheres immersed in a polarizable medium (see Fig.1) and separated by a centre-to-centre separation $h$. For sphere $i\in\left\{1,2\right\}$ with a radius $a$, its free charge $q_{i}$ is concentrated at the particle’s centre $\mathbf{z}_{i}=\left[0,0,z_{i}\right]$, and its bound charge spans the particle’s surface. At any position $\mathbf{r}$, the volume density $\rho(\mathbf{r})$ of total charge is a sum of the volume density $\rho_{\text{f}}(\mathbf{r})$ of free charge and the volume density $\rho_{\text{b}}(\mathbf{r})$ of polarization-induced bound charge:
@@
where the electrostatic potential $\phi(\boldsymbol{r})$ is a sum of Coulomb contributions from all free charge and bound charge elsewhere in the system. This potential approaches zero at any position far away from the pair of dielectric spheres. According to Gauss’ law, we have
@@
where $\epsilon_{\text{r}}(\mathbf{r})$ is the relative permittivity and $\epsilon_{0}$ the permittivity of free space. Since the polarizable medium and the two dielectric spheres are each homogeneous, the space-dependent relative permittivity is given by
@@ II.2 Computational method
respectively, where $Y_{\ell}^{m}(\theta_{i},\psi_{i})$ denotes the spherical harmonics function with degree $\ell$ and order $m$. The expansion coefficients and $A_{i,\ell m}$ and $B_{i,\ell m}$ were evaluated numerically via the aforementioned hybrid computational method  Gan et al. (2016, 2019), and truncated up to some prescribed order $p$ to achieve a desired accuracy. The electrostatic energy can then be computed using Eqs. (3) and (10). The electrostatic force on sphere $i$ is then given by
@@
according to Eqs. (9) and (10) and the corresponding interfacial boundary condition. For the sphere’s effective dipole Lindgren et al. (2018), the positive charge
@@
as well as the negative charge
@@
respectively. For sphere $i$, the absolute quantity $\widetilde{q}_{i}\coloneqq|\widetilde{q}_{i}^{\pm}|$ is the magnitude of induced charge, and the positions $\widetilde{z}_{i}^{\pm}$ are the charge-weighted average positions of the two types of induced charge. Due to azimuthal symmetry, the induced charges are located on the $z$-axis and thus the $xy$-coordinates can be omitted.
III Results and Discussion
To investigate how the electrostatic interaction between two charged dielectric spheres is influenced by the polarizability of the medium, we first consider the scenario of a less polarizable medium for both spheres (scenario 1: $k_{1}>k_{0}$ and $k_{2}>k_{0}$), and then the scenario of a more polarizable medium for both spheres (scenario 2: $k_{1}<k_{0}$ and $k_{2}<k_{0}$). For each scenario, we consider cases of like-charge interactions as well as cases of opposite-charge interactions, and look into how the nature of interparticle interaction and the physical picture of charge separation vary as the particles approach each other. In all cases, we consider combinations of values of $q_{i}$ and $k_{i}$ that correspond to a dominant mechanism of polarization of sphere 1 on the left by sphere 2 on the right (Fig. 1), and look into how our results vary with the strength of such polarization. For scenario 1, the interaction between two like-charged spheres could be repulsive or attractive but the interaction between two oppositely charged spheres must be attractive, because any polarization-induced interaction must be attractive. Conversely, for scenario 2, the interaction between two like-charged spheres must be repulsive but the interaction between two oppositely charged spheres could be attractive or repulsive, because any polarization-induced interaction must be repulsive.
Scenario 1: A less polarizable medium for both spheres

### v2

@@ Abstract
The mechanisms of electrostatic interactions between two charged dielectric spheres inside a polarizable medium have been investigated, in terms of hypothetical effective dipoles that depict how the positive and negative charge in each particle are separated. Our findings, which revealed that it is possible for polarization-induced opposite-charge repulsion to occur at short interparticle separations if the dielectric constant of the medium is greater than the dielectric constants of both spheres, provide insights into the physics of charge separation in each sphere and of polarization in the medium behind such counterintuitive behaviour.
@@ I Introduction
While the classic Coulomb’s law Jackson (2021) describes the electrostatic force between two point charges in a vacuum, it does not necessarily describe the electrostatic force between two charged dielectric particles. In some cases, there could be a significant deviation from the Coulomb force at small separations, resulting in the occurrence of polarization-driven like-charge attraction Feng (2000); Bichoutskaia et al. (2010); Xu (2013); Freed (2014); Qin et al. (2016, 2016); Lindgren et al. (2018); Qin (2019); Chan (2020); Li et al. (2024); Gorman et al. (2024). On the other hand, charged particles immersed in an electrolyte could also exhibit anti-Coulomb behaviour, including cases of like-charge attraction  Allahyarov et al. (1998); Levin et al. (1999); Linse and Lobaskin (1999); Moreira and Netz (2001); Nagornyak et al. (2009); Zhao et al. (2016); Buyukdagli (2017); Dos Santos and Levin (2019); Wills et al. (2024); Wang et al. (2024) and also cases of opposite-charge repulsion Aranda-Espinoza et al. (1999); Zhang and Shklovskii (2004); Leunissen et al. (2005); Wills et al. (2024), but such behaviour is not purely driven by the polarizability of the particles nor that of the surrounding medium. Such cases, where ionic effects are significant, could be investigated theoretically via mean-field approaches Budkov and Kolesnikov (2022); Kolesnikov et al. (2022), the variational-field theory Budkov and Brandyshev (2023), the statistical-field theory Brandyshev and Budkov (2023), the self-consistent field theory Budkov and Kalikin (2023), or a thermomechanical approach Budkov et al. (2024). While it is physically impossible for two dielectric particles in a vacuum to exhibit opposite-charge repulsion, by taking into account the polarizability of the medium it would be worthwhile to investigate whether opposite-charge repulsion could also occur as a pure consequence of polarization effects, as in the case of like-charge attraction between two dielectric particles in a vacuum.
@@
To understand the physical mechanisms behind the occurrence of polarization-driven like-charge attraction between two polarizable particles, much research has been conducted by means of 1) accurate evaluations of the interparticle force or energy Bichoutskaia et al. (2010); Khachatourian et al. (2014); Lindgren et al. (2016, 2018); Lian and Qin (2018); Li et al. (2024) that take into account all possible higher-order interactions, or 2) effective-dipole-based modelsChan (2020) or computationsLindgren et al. (2018); Li et al. (2024) that provide a physical picture of charge separation in each particle. For the former, a multiple scattering formalism Freed (2014); Qin et al. (2016, 2016); Qin (2019) has been developed, where analytic solutions for generally many-body systems could be obtained through an application of the perturbative many-body expansion to the interaction energy. On the other hand, there have been various attempts to develop efficient numerical methods for accurate computations of force or energy: conventional approaches, such as the finite element method Feng (2000) and the boundary element method Gorman et al. (2024); Ruan et al. (2022); Hassan and Stamm (2021); Bramas et al. (2021), typically involve the discretization of a volume or surface. At small interparticle separations, such approaches become computationally expensive, because a high resolution of the discretization is required. For this reason, some more efficient approaches have been developed. These include 1) the image charge method  Wang and Wang (2013); Xu (2013) where, for each particle, the polarization field is approximated as a field generated by the image charges in the other particle and by the multiply-reflected image charges due to the presence of two interfaces, 2) the method of moments  Bichoutskaia et al. (2010); Derbenev et al. (2016); Lindgren et al. (2016); Siryk et al. (2021) where the polarization-related potential is described in terms of a multipole expansion, the coefficients of which are determined by substituting the expansion into the Poisson equation and the corresponding boundary conditions, and 3) a hybrid computational method that employs the image charge method to compute the fast varying near-field interactions and the method of moments to solve the remaining smooth far-field interactions.
@@
In this research, we used a combination of the hybrid computational method with the effective-dipole approach to investigate the possibility of polarization-induced opposite-charge repulsion between two dielectric spheres in close proximity inside a polarizable medium. This effective-dipole approach averages over not only the first-order dipole contributions but also all quadrupole and higher-order contributions. Our findings revealed that such a possibility exists if the dielectric constant of the medium is greater than the dielectric constants of both spheres, in which case the repulsive interaction corresponds to an anti-parallel alignment of the particles’ effective dipoles Lindgren et al. (2018) and hence a dominant like-charge repulsion between the polarization charges of the particles. The problem of how the electrostatic interactions of charged dielectric particles can be mediated by the polarizability of the medium is of fundamental interest and importance to the scientific community. In this paper, we provide a unified understanding of how the counterintuitive polarization-driven phenomena of like-charge attraction and opposite-charge repulsion can be triggered, where our effective-dipole analysis provides insights into the physics of charge separation in each particle under the influence of a polarizable medium.
@@
This paper is organized as follows. In Sec. II, we present a theoretical model of two charged dielectric spheres, on which our research is based, along with a brief review of the hybrid computational method Gan et al. (2019) that we employed to compute the interparticle force and the effective-dipole moment of each particle  Lindgren et al. (2018). In Sec. III, we present our findings for four different scenarios of interparticle interactions. In Sec. IV, we discuss the mechanisms of interparticle interactions and the corresponding orientations and relative positions of effective dipoles for each scenario. In Sec. V, we summarize our findings and discuss their implications.
@@ II.1 Model of dielectric spheres in a polarizable medium
Figure 1: Schematic illustration of two dielectric spheres, immersed in a medium of dielectric constant $k_{0}$ and separated by a centre-to-centre separation $h$. For sphere $i\in\left\{1,2\right\}$ with a radius $a_{i}$ and a dielectric constant $k_{i}$, its free charge $q_{i}$ is concentrated at the particle’s centre $\mathbf{z}_{i}$, and its bound charge spans the particle’s surface.
Consider two dielectric spheres immersed in a polarizable medium (see Fig.1) and separated by a centre-to-centre separation $h$. For sphere $i\in\left\{1,2\right\}$ with a radius $a_{i}$, its free charge $q_{i}$ is concentrated at the particle’s centre $\mathbf{z}_{i}=\left[0,0,z_{i}\right]$, and its bound charge spans the particle’s surface. At any position $\mathbf{r}$, the volume density $\rho(\mathbf{r})$ of total charge is a sum of the volume density $\rho_{\text{f}}(\mathbf{r})$ of free charge and the volume density $\rho_{\text{b}}(\mathbf{r})$ of polarization-induced bound charge:
@@
where the electrostatic potential $\phi(\boldsymbol{r})$ is a sum of Coulomb contributions from all free charge and bound charge elsewhere in the system. It is worth pointing out that the electrostatic energy $U$ duly excludes the interaction between any given point charge and the potential it generates, where such self-interaction would inappropriately make the energy divergent. This potential approaches zero at any position far away from the pair of dielectric spheres. According to Gauss’ law, we have
@@
where $\epsilon_{\text{r}}(\mathbf{r})$ is the relative permittivity and $\epsilon_{0}$ the permittivity of free space. Since the polarizable medium and the two dielectric spheres are each homogeneous, the position-dependent relative permittivity is given by
@@ II.2 Computational method
respectively, where $Y_{\ell}^{m}(\theta_{i},\psi_{i})$ denotes the spherical harmonics function with degree $\ell$ and order $m$. The expansion coefficients and $A_{i,\ell m}$ and $B_{i,\ell m}$ were evaluated numerically via the aforementioned hybrid computational method  Gan et al. (2016, 2019), and truncated up to some prescribed order $p$ to achieve a desired accuracy. Although we do not have general analytic solutions for the expansion coefficients $A_{i,\ell m}$ and $B_{i,\ell m}$, there is one special case where the coefficient can be determined analytically and is found to be $h$-independent  Gan et al. (2019): $A_{i,00}=q_{i}$. Any $h$-dependence of monopole strength arises from a screening effect of mobile ions, which is not considered in our study of purely polarization-driven interparticle interactions. Therefore the model considered in our study corresponds to the limiting case of zero ionic strength Fisher et al. (1994). Upon a computation of the electrostatic energy $U$ using Eqs. (3) and (10), the electrostatic force on sphere $i$ can be evaluated as follows:
@@
according to Eqs. (9) and (10) and the corresponding interfacial boundary condition. For the sphere’s effective dipole Lindgren et al. (2018), the total positive polarization-charge
@@
as well as the total negative polarization-charge
@@
respectively. Any polarization effect, which by definition corresponds to a spatial separation of positive and negative charge, does not contribute to the net charge of a sphere, and therefore the total polarization charge of a sphere must be zero, i.e.
$\widetilde{q}_{i}^{+}+\widetilde{q}_{i}^{-}=0\;\$ (19)
@@
Eq. (19) has been verified in all of our numerical simulations. For sphere $i$, the absolute quantity $\widetilde{q}_{i}\coloneqq|\widetilde{q}_{i}^{\pm}|$ is the magnitude of induced charge, and the positions $\widetilde{\mathbf{r}}_{i}^{\pm}$ are the charge-weighted average positions of the two types of induced charge. Due to azimuthal symmetry, the induced charges are all located on the $z$-axis such that the positions $\widetilde{\mathbf{r}}_{i}^{\pm}$ are given by $\widetilde{\mathbf{r}}_{i}^{+}\coloneqq[0,0,\widetilde{z}_{i}^{+}]$ and $\widetilde{\mathbf{r}}_{i}^{-}\coloneqq[0,0,\widetilde{z}_{i}^{-}]$, respectively.
III Scenarios of interparticle interactions
Figure 2: Rescaled electrostatic force and rescaled polarization force on sphere 2 for a pair of equal-sized dielectric spheres inside a polarizable medium: a) $k_{1}/k_{0}=0.001$, $k_{2}/k_{0}=0.08$ and $q_{2}/q_{1}=8$ (scenario A); (b) $k_{1}/k_{0}=30$, $k_{2}/k_{0}=4$ and $q_{2}/q_{1}=7$ (scenario B); (c) $k_{1}/k_{0}=0.001$, $k_{2}/k_{0}=0.08$ and $q_{2}/q_{1}=-8$ (scenario C); (d) $k_{1}/k_{0}=30$, $k_{2}/k_{0}=4$ and $q_{2}/q_{1}=-7$ (scenario D).
Figure 3: Rescaled electrostatic force and rescaled polarization force on sphere 2 for a pair of unequal-sized dielectric spheres at $|q_{1}|=|q_{2}|$, $a_{2}/a_{1}=0.2$ and $k_{1}/k_{0}=k_{2}/k_{0}=k$ inside a polarizable medium: (a) $q_{1}=q_{2}$ and $k=0.25$ (scenario A); (b) $q_{1}=q_{2}$ and $k=4$ (scenario B); (c) $q_{1}=-q_{2}$ and $k=0.25$ (scenario C); (d) $q_{1}=-q_{2}$ and $k=4$ (scenario D).
Figure 4: Four scenarios of electrostatic interactions under consideration: (a) $k_{1}<k_{0}$, $k_{2}<k_{0}$ and $q_{2}/q_{1}>0$ (scenario A); (b) $k_{1}>k_{0}$, $k_{2}>k_{0}$ and $q_{2}/q_{1}>0$ (scenario B); (c) $k_{1}<k_{0}$, $k_{2}<k_{0}$ and $q_{2}/q_{1}<0$ (scenario C); (d) $k_{1}>k_{0}$, $k_{2}>k_{0}$ and $q_{2}/q_{1}<0$ (scenario D).
To investigate how the electrostatic interaction between two charged dielectric spheres depends on the dielectric constants $k_{i}$, free charges $q_{i}$ and radii $a_{i}$ of the spheres ($i=1,2$) and on the dielectric constant $k_{0}$ of the polarizable medium, we first consider two types of scenarios: (i) two equal-sized spheres with different dielectric constants $k_{i}$ and charge magnitudes $|q_{i}|$, and (ii) two unequal-sized spheres sharing the same dielectric constant $k_{i}$ and charge magnitude $|q_{i}|$. For each type of scenarios, we consider how the total electrostatic force and polarization force vary with the interparticle separation at four possible combinations of charge ratio ($q_{2}/q_{1}>0$ for like-charge interactions, or $q_{2}/q_{1}<0$ for opposite-charge interactions) and medium polarizability ($k_{i}<k_{0}$ for a more polarizable medium, or $k_{i}>k_{0}$ for a less polarizable medium): (a) $q_{2}/q_{1}>0$ and $k_{i}<k_{0}$ (scenario A); (b) $q_{2}/q_{1}>0$ and $k_{i}>k_{0}$ (scenario B); (c) $q_{2}/q_{1}<0$ and $k_{i}<k_{0}$ (scenario C); (d) $q_{2}/q_{1}<0$ and $k_{i}>k_{0}$ (scenario D). Without loss of generality, we adopted the parameter values $k_{0}=1$, $q_{1}=1\text{e}$ and $a_{1}=1\mu\text{m}$ in our numerical simulations.
Fig. 2 shows how the rescaled electrostatic force and rescaled polarization force vary with the interparticle separation for cases of equal-sized spheres with different dielectric constants $k_{i}$ and charge magnitudes $|q_{i}|$. For scenarios A and C, where the dielectric constants of the spheres are smaller than that of the polarizable medium, the polarization force is all the way repulsive, regardless of whether the spheres exhibit like-charge or opposite-charge interactions. For scenario A, the Coulomb force is inherently repulsive, and the polarization force simply enhances the repulsive nature of the interparticle interaction. For scenario C, the Coulomb force is inherently attractive, and the polarization force results in a switch of the total electrostatic force from attraction to repulsion for decreasing separation. For scenarios B and D, where the dielectric constants of the spheres are larger than that of the polarizable medium, the polarization force is all the way attractive. For scenario B, the Coulomb force is inherently repulsive, and the polarization force results in a switch of the total electrostatic force from repulsion to attraction for decreasing separation. For scenario D, the Coulomb force is inherently attractive, and the polarization force simply enhances the attractive nature of the interparticle interaction. For all cases, at large interparticle separations, the polarization force approaches zero and the total electrostatic force becomes practically equal to the Coulomb force. Fig. 3 shows corresponding results for cases of unequal-sized spheres sharing the same dielectric constant $k_{i}$ and charge magnitude $|q_{i}|$. The results, which display the same qualitative features as those shown in Fig. 2, suggest a universal picture of electrostatic interactions upon the dominant polarization of one sphere by the other, as summarized in Fig. 4:
If the dielectric constant of the surrounding medium exceeds the dielectric constants of both spheres, i.e. $k_{1}<k_{0}$ and $k_{2}<k_{0}$, the polarization within the system results in an additional repulsive interaction that either strengthens the Coulomb repulsion between two like-charged particles ($q_{2}/q_{1}>0$) in scenario A or weakens (or even overcomes) the Coulomb attraction between two oppositely charged particles ($q_{2}/q_{1}<0$) in scenario C. Else if the dielectric constant of the surrounding medium is smaller than the dielectric constants of both spheres, i.e. $k_{1}>k_{0}$ and $k_{2}>k_{0}$, the polarization within the system results in an additional attractive interaction that either weakens (or even overcomes) the Coulomb repulsion between two like-charged particles ($q_{2}/q_{1}>0$) in scenario B or strengthens the Coulomb attraction between two oppositely-charged particles ($q_{2}/q_{1}<0$) in scenario D.
IV Mechanisms of interparticle interactions
Without loss of generality, we have conducted an effective-dipole analysis of the mechanisms of interparticle interactions for cases of two equal-sized spheres with different dielectric constants $k_{i}$ and charge magnitudes $|q_{i}|$, with a focus on how the nature of interaction varies as the spheres approach each other. For each scenario, we consider various combinations of $q_{i}$ and $k_{i}$ that correspond to a dominant mechanism of polarization of sphere 1 on the left by sphere 2 on the right (Fig. 1), and look into how our results vary with the strength of such polarization for decreasing interparticle separation.
Cases of a less polarizable medium